a ball is thrown from a 128ft high balcany with the initial velocity of 32ft / sec the height of the ball can be molded by the function h (t) = -t squared +32t +128

when will the ball be at the highest point amd how high is that point how long will it take to reach the ground?


1
Expert's answer
2020-01-20T05:22:34-0500

At the highest point:

"v(t)=-2t+32=0"

"t=16\ s"

"H=-16^2+32(16)+128=384\ ft"

The time to reach the ground:


"h=0\to -t^2 +32t +128=0"

"T=35.6\ s"